4. Calculate the enthalpy of the combustion reaction of methanol using the binding energies.


5. In a calorimeter, a student dissolved 4.25 g of solid ammonium nitrate in 60.0 ml of water with an initial temperature of 22.0 °C. When the ammonium nitrate was completely dissolved, the temperature of the solution was 16.9 °C. Based on the measurement results, calculate the heat of dissolution of ammonium nitrate in kJ/mol.


1
Expert's answer
2021-02-02T01:24:47-0500

Q159073


4. Calculate the enthalpy of the combustion reaction of methanol using the binding energies.


Solution:


The combustion of methanol is given as


2CH3OH(g) + 3 O2 (g) ==> 2 CO2 (g) + 4 H2O (g)





The calculated heat of combustion will vary depending on the data provided to us.

I have used the data given in the table and calculated

The heat of combustion of methanol = -1340 kJ


In the reaction the coefficient of CH3 OH is 2.

So enthalpy of combustion for 1 mole of methanol will be = -1340 kJ/2 mol = -670kJ/mol .


Enthalpy for combustion reaction will be = -1340 kJ


and enthalpy of reaction per mole of methanol will be - 670 kJ/mol.





5. In a calorimeter, a student dissolved 4.25 g of solid ammonium nitrate in 60.0 ml of water with an initial temperature of 22.0 °C. When the ammonium nitrate was completely dissolved, the temperature of the solution was 16.9 °C. Based on the measurement results, calculate the heat of dissolution of ammonium nitrate in kJ/mol.

Solution :


Let us assume that the density of water is 1.00g/mL

So the mass of 60.0 mL of water will be equal = 60.0 grams.


4.25g of NH4NO3 is added to 60 grams of water.

So mass of solution = 4.25g + 60.0 g = 64.25 grams.


Step 1: To find the heat given by the dissolution of ammonium nitrate.


We will use the specific heat capacity formula


Q = m * s *  ∆T = 64.25 g * 4.184 J/g 0C * ( 16.9 0C - 22.0 0C)

Q = 64.25 * 4.184 J/g 0C * (-5.1 0C )

Q = -1370.99 Joules.


Step 2: To find Heat of dissolution of ammonium nitrate in kJ/mol


grams of NH4NO3 = 4.25 grams


molar mass of NH4NO3 = 2 * atomic mass of N + 4 * atomic mass of H + 3 * atomic mass of O

= 2 * 14.007 g/mol + 4 * 1.00794 g/mol + 3 * 15.999 g/mol

= 28.014g/mol + 4.03176 g/mol +47.997 g/mol

= 80.043g/mol



moles of NH4NO3 = 4.25 g NH4NO3 * 1mol NH4NO3 / 80.043 g NH4 NO3


= 0.05309 moles



Heat of dissolution of ammonium nitrate = -1370.99 Joules / 0.05309 mol


= -25820 Joules/mol


Convert this to kilojoules using 1kilojoule = 1000 Joule



in kilojoule the answer will be = -25820 J/mol * 1kJ/ 1000 J

= -258 kilojoules per mole



The heat of dissolution of ammonium nitrate is -258 kilojoules per mole